# Convergence of Decentralized Stochastic Subgradient-based Methods for Nonsmooth Nonconvex functions

### v3

@@ Abstract
Keywords: Nonsmooth optimization, decentralized optimization, stochastic subgradient-based method, random reshuffling, with-replacement sampling, conservative field, Lyapunov function.
@@ 1 Introduction
Here, for each $i\in[d]:=\{1,2,\ldots,d\}$, ${\bm{x}}_{i}$ refers to the local variable of agent $i$. Moreover, the set of nodes $\mathtt{V}=[d]:=\{1,\ldots,d\}$ represents the set of agents, while the set of edges $\mathtt{E}\subseteq\mathtt{V}\times\mathtt{V}$ represents the set of communication links. An edge $(i,j)\in\mathtt{E}$ refers to that agents $i$ and $j$ are neighbors and can communicate with each other. Furthermore, for any $i\in[d]$, $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz continuous (possibly nonconvex and nonsmooth) cost function exclusively known to the agent $i$ and takes the empirical-risk formulation [1]. More precisely, for any $i\in[d]$, there exists $F_{i}:\mathbb{R}^{n}\times\Omega\to\mathbb{R}$ such that
@@
where ${\bm{s}}_{l}$ is a random data vector supported on a local finite set $\mathcal{S}_{i}$. In real-world applications, the size of $\mathcal{S}_{i}$ is often sufficiently large, in the sense that computing the exact (sub)gradient of $f_{i}$ is very costly, which makes stochastic (sub)gradient-based methods inevitable. Problem (DOP) has wide applications in machine learning [2, 3], signal processing [4, 5] and control [6], etc.
@@ 1.1 Existing works on decentralized stochastic optimization
Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [7, 8, 9] is one of the most fundamental methods for solving (DOP). In DSGD, each agent locally averages its model parameters with those of its neighbors and updates along the direction of the negative local stochastic gradient, which can be expressed as
@@
Based on DSGD, a wide range of variants have been developed for solving (DOP). Some works introduce auxiliary variables within DSGD to track the global stochastic gradient of the objective function, such as GNSD [10] and DSGT [11]. To accelerate convergence, some recent works [12, 13] propose the decentralized momentum SGD by integrating the heavy-ball momentum [14] into DSGD. Moreover, [15] combines the stochastic gradient tracking method with momentum acceleration. In addition, [16, 17] introduce random reshuffling (RR) [18] sampling techniques (i.e., randomly samples an index $i_{k}$ of $[N]$ without replacement at each step $k$ in agent $i$, ensuring that each index appears exactly once per epoch.) into DSGD and DSGT, and provide theoretical evidence that RR outperforms with-replacement sampling. A series of studies [9, 19, 20, 21, 22] have established theoretical guarantees for these decentralized methods with continuously differentiable nonconvex objective functions.
@@
Much attention has also been directed towards decentralized nonsmooth optimization. Some existing works, such as ProxDGD [23], NEXT [24], and PG-EXTRA [25] develop decentralized proximal gradient type methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus is Clarke regular. Several existing works [26, 27, 28, 29] extend the convergence analysis of Prox-DGD and NEXT from the deterministic case to the stochastic case, under additional assumptions regarding boundedness and heterogeneity. Additionally, [30] proposes stoDPSM for $\rho$-weakly convex objectives and shows the global convergence while employing the Moreau envelope for measuring the stationarity. [31] improves the convergence analysis for DSGT and proves its convergence for weakly convex objectives. From the view of distributed gradient flow, [32] establishes the global convergence of DSGD for Clarke regular function with decreasing consensus parameter. The detailed comparisons between the existing results on decentralized SGD-based methods in nonsmooth nonconvex optimization are listed in Table 1.
@@
Despite achieving nice theoretical results, all the aforementioned works are only capable of handling nonsmooth but (Clarke) regular [33] objective functions in (DOP), while some of these works evaluate the efficiency of their developed methods in the decentralized training of neural networks. However, nonsmooth activation functions, including ReLU and leaky ReLU, have served as popular building blocks for modeling neural networks, resulting in loss functions that are nonsmooth and lack Clarke regularity. To the best of our knowledge, [34] is the only recent work that investigates the convergence of decentralized SGD-based method without requiring Clarke regularity. They propose a decentralized first-order and zero-order online algorithm called Me-DOL, but it requires several strong assumptions, such as global Lipschitz function and $L^{2}$ bounded stochastic subgradients. As a result, how to establish convergence properties for decentralized stochastic subgradient-based methods in nonsmooth, nonconvex settings remains an area for further exploration.
@@
stoDPSM [30] GD Square summable, $\eta_{k+1}/\eta_{k}\to 1$ $\rho$-weakly convex, bounded subgradient WRS, bounded second moment
S-NEXT [26] GT Square summable Proximal, weakly convex, bounded subgradient WRS, bounded noise
ProxGT [28] GT Constant Proximal, weakly convex, mean-squared smoothness WRS, bounded variance
DGF [32] GD $\eta_{k}=O(k^{-p}),p\in(0,1]$ Clarke regular, weakly coercive, decreasing consensus parameter WRS, bounded variance
ME-DOL [34] GD $\eta_{k}=\Theta(\frac{1}{\sqrt{K}}),k\in[K]$ Global Lipschitz continuous WRS, bounded variance, bounded second moment
@@
Our work (DSM) GD/GT /GD-M diminishing $\eta_{k}$ Path-differentiable, Coercive RR
Table 1: A brief comparison of existing decentralized SGD-based methods for nonsmooth nonconvex optimization. Here, “GD”, “GT” and “GD-M” are abbreviations of “gradient descent”, “gradient tracking” and “gradient descent with momentum technique”; “WRS” and “RR” are abbreviations of “with-replacement sampling” and “random reshuffling”; The term “conditionally controlled noise” means that the evaluation noise is bounded conditional on $\mathcal{F}_{k}=\sigma(\{{\bm{x}}_{i,l},l\leq k,i\in[d]\})$. The concept ”path-differentiable” is defined in Definition 2.6.
@@ 1.2 Existing works on stochastic subgradient-based methods for nonsmooth nonconvex optimization
In the training of nonsmooth neural networks, one major challenge comes from computing the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, such as TensorFlow, PyTorch, and JAX, automatic differentiation (AD) algorithms have become the default choice for computing subdifferential. Utilizing the chain rule, AD algorithms recursively construct the so-called “subgradients” through the composition of Jacobians of each block of the neural network, yet they neglect the underlying nonsmoothness. As the chain rule fails for loss functions lacking Clarke regularity, the so-called “subgradient” may not necessarily be contained in its Clarke subdifferential (see examples in [35]). This limitation consequently renders the theoretical analyses of existing literature inapplicable to such functions.
@@
To characterize the behavior of AD algorithms, [36] introduces the concept of path-differentiable functions, which constitute a broad function class, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [37]. More importantly, the set of path-differentiable functions is general enough to cover almost all loss functions in neural network tasks, encompassing most of the loss functions without Clarke regularity. The study [36] further demonstrates that when applied to path-differentiable loss functions, the outputs of AD algorithms are contained within a conservative field, which is a generalization of the Clarke subdifferential. Conservative field admits chain rules for path-differentiable functions, thereby preserving several good properties. For a comprehensive discussion of this concept, readers are directed to Section 2.3.
@@
Based on the concepts of path-differentiable functions and conservative field, some recent works [36, 38, 39, 40, 41, 42, 43, 44] leverage the ordinary differential equation (ODE) approach [45, 46, 47, 48] to study the behavior of stochastic subgradient-based methods in the non-distributed setting. Although these studies have achieved significant progress in establishing convergence properties for stochastic subgradient-based methods, the theoretical results of many approaches cannot avoid infinitely many spurious critical points due to the gap between the conservative field and the Clarke subdifferential [49, 50]. Moreover, extending existing results to the multi-agent scenario is non-trivial, as the update schemes of all the agents are coupled through local communication under decentralized settings. Given the convergence properties that need improvement and limited existing results in the multi-agent scenario, developing a unified framework for establishing convergence properties for decentralized stochastic subgradient-based methods is of great importance and worth exploring.
@@ 1.3 A general framework for decentralized stochastic subgradient-based methods
${\bm{X}}_{k}:=$ $\left[{\bm{x}}_{1,k},{\bm{x}}_{2,k},\ldots,{\bm{x}}_{d,k}\right],$ (1.3)
$\Xi_{k+1}:=$ $\left[\xi_{1,k+1},\xi_{2,k+1},\ldots,\xi_{d,k+1}\right],$
${\bm{H}}_{k}:=$ $[\nabla f_{1}({\bm{x}}_{1,k}),\ldots,\nabla f_{d}({\bm{x}}_{d,k})].$
@@ 1.4 Contributions
We propose a unified framework (DSM) that encompasses a variety of decentralized stochastic subgradient-based methods. We verify that decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T), and DSGD with momentum (DSGD-M) all fit into the framework (DSM). In particular, a decentralized variant of SignSGD [51], denoted as DSignSGD, is also enclosed by (DSM).
@@
Based on our theoretical results of the proposed framework (DSM), we prove that DSGD, DSGD-M, DSGD-T and DSignSGD achieve global convergence in the random reshuffling case and high-probability global convergence in the with-replacement sampling case. We also show that with random initialization, DSGD, DSGD-M and DSignSGD can avoid the spurious critical points in the sense of conservative field in minimizing path-differentiable functions. Although these algorithms are widely used in the training of nonsmooth neural networks, existing theoretical results are limited to Clarke regular objective functions. Our theoretical findings first fill the gap in solving Problem (DOP) with objectives lacking Clarke regularity.
@@ 2.1 Notations
[…]
@@
Moreover, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{1}}_{d}\in\mathbb{R}^{d}$ represent a vector of all $1$’s, ${\bm{e}}_{i}\in\mathbb{R}^{d}$ represent $[0,\ldots,1,\ldots,0]^{\top}$, where $1$ is the $i$-th component. For two integers $i$ and $j$, $i\land j$ represents $\min\{i,j\}$.

### v4

@@ Abstract
Keywords: Nonsmooth optimization, decentralized stochastic subgradient-based method, random reshuffling, with-replacement sampling, conservative field, Lyapunov function.
@@ 1 Introduction
Here, for each $i\in[d]:=\{1,2,\ldots,d\}$, ${\bm{x}}_{i}$ refers to the local variable of agent $i$. Moreover, the set of nodes $\mathtt{V}=[d]:=\{1,\ldots,d\}$ represents the set of agents, while the set of edges $\mathtt{E}\subseteq\mathtt{V}\times\mathtt{V}$ represents the set of communication links. An edge $(i,j)\in\mathtt{E}$ refers to that agents $i$ and $j$ are neighbors and can communicate with each other. Furthermore, for any $i\in[d]$, $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz continuous (possibly nonconvex and nonsmooth) cost function exclusively known to the agent $i$ and takes the empirical-risk formulation [shapiro2021lectures]. More precisely, for any $i\in[d]$, there exists $F_{i}:\mathbb{R}^{n}\times\Omega\to\mathbb{R}$ such that
@@
where ${\bm{s}}_{l}$ is a random data vector supported on a local finite set $\mathcal{S}_{i}$. In real-world applications, the size of $\mathcal{S}_{i}$ is often sufficiently large, in the sense that computing the exact (sub)gradient of $f_{i}$ is very costly, which makes stochastic (sub)gradient-based methods inevitable. Problem (DOP) has wide applications in machine learning [jain2017non, bottou2018optimization], signal processing [mateos2012distributed, cohen2016distributed] and control [bolognani2014distributed], etc.
@@ 1.1 Existing works on decentralized stochastic optimization
Motivated by stochastic gradient descent (SGD), the decentralized stochastic gradient descent (DSGD) [nedic2009distributed, jiang2017collaborative, lian2017can] is one of the most fundamental methods for solving (DOP). In DSGD, each agent locally averages its model parameters with those of its neighbors and updates along the direction of the negative local stochastic gradient, which can be expressed as
@@
Based on DSGD, a wide range of variants have been developed for solving (DOP). Some works introduce auxiliary variables within DSGD to track the global stochastic gradient of the objective function, such as GNSD [lu2019gnsd] and DSGT [pu2021distributed]. To accelerate convergence, some recent works [yu2019linear, gao2020periodic] propose the decentralized momentum SGD by integrating the heavy-ball momentum [polyak1964some] into DSGD. Moreover, [gao2023distributed] combines the stochastic gradient tracking method with momentum acceleration, [chen2023convergence] proposes a general framework for decentralized adaptive methods such as Adam, AdaGrad, and AMSGrad, and show that under suitable conditions, for nonconvex smooth objectives, if the original adaptive method converges, then its decentralized counterpart also converges. In addition, [huang2023distributed, huang2023distributed2] introduce random reshuffling (RR) [de2020random] sampling techniques (i.e., randomly samples an index $i_{k}$ of $[N]$ without replacement at each step $k$ in agent $i$, ensuring that each index appears exactly once per epoch.) into DSGD and DSGT, and provide theoretical evidence that RR outperforms with-replacement sampling. A series of studies [lian2017can, george2019distributed, yuan2021decentlam, zhang2019decentralized, xin2021improved] have established theoretical guarantees for above decentralized SGD-type methods with continuously differentiable nonconvex objective functions. Recently, a line of works [Yan2023CompressedDP, Bylinkin2024AcceleratedMW, condat2024locodl, xu2023parallel, sun2024role, Liu2025CompressedDM] have focused on developing distributed SGD-type methods that reduce communication and handle asynchrony and heterogeneity. For instance, [condat2024locodl] couples local training with communication compression to obtain a communication-efficient distributed optimization method with theoretical guarantees with strongly convex objectives. [xu2023parallel] analyzes asynchronous adaptive stochastic gradient methods, showing that AMSGrad-type schemes can be parallelized under bounded staleness while retaining convergence guarantees with smooth objective functions. From the perspective of federated learning, [sun2024role] studies the role of server momentum and develop a more general server-momentum framework, which highlights how momentum can mitigate the adverse effects of heterogeneous data and system asynchrony with smooth objective functions.
@@
Much attention has also been directed towards decentralized nonsmooth optimization. Some existing works, such as ProxDGD [zeng2018nonconvex], NEXT [di2016next], and PG-EXTRA [shi2015proximal] develop decentralized proximal gradient type methods for (DOP), under the assumption that $f_{i}$ is composed of a differentiable term $g_{i}$ and a convex regularization term $r_{i}$ for all $i\in[d]$. In this context, the objective function is weakly convex and thus is Clarke regular. Several existing works [di2019distributed, niu2021distributed, xin2021stochastic, xiao2023one, zhou2025decentralized] extend the convergence analysis of Prox-DGD and NEXT from the deterministic case to the stochastic case, under additional assumptions regarding boundedness and heterogeneity. Additionally, [chen2021distributed] proposes stoDPSM for $\rho$-weakly convex objectives and shows the global convergence while employing the Moreau envelope for measuring the stationarity. [koloskova2021improved] improves the convergence analysis for DSGT and proves its convergence for weakly convex objectives. From the view of distributed gradient flow, [swenson2022distributed] establishes the global convergence of DSGD for Clarke regular function with decreasing consensus parameter. The detailed comparisons between the existing results on decentralized SGD-based methods in nonsmooth nonconvex optimization are listed in Table 1.
@@
Despite achieving nice theoretical results, all the aforementioned works are only capable of handling nonsmooth but (Clarke) regular [clarke1990optimization] objective functions in (DOP), while some of these works evaluate the efficiency of their developed methods in the decentralized training of neural networks. However, nonsmooth activation functions, including ReLU and leaky ReLU, have served as popular building blocks for modeling neural networks, resulting in loss functions that are nonsmooth and lack Clarke regularity. To the best of our knowledge, [sahinoglu2024online] is the only recent work that investigates the convergence of decentralized SGD-based method without requiring Clarke regularity. They propose a decentralized first-order and zero-order online algorithm called Me-DOL, but it requires several strong assumptions, such as global Lipschitz function and $L^{2}$ bounded stochastic subgradients. As a result, how to establish convergence properties for decentralized stochastic subgradient-based methods in nonsmooth, nonconvex settings remains an area for further exploration.
@@
stoDPSM [chen2021distributed] GD Square summable, $\eta_{k+1}/\eta_{k}\to 1$ $\rho$-weakly convex, bounded subgradient WRS, bounded second moment
S-NEXT [di2019distributed] GT Square summable Proximal, weakly convex, bounded subgradient WRS, bounded noise
ProxGT [xin2021stochastic] GT Constant Proximal, weakly convex, mean-squared smoothness WRS, bounded variance
Prox-DASA [xiao2023one] GD/GT $\eta_{k}=\Theta(\frac{1}{\sqrt{K}}),k\in[K]$ Proximal, weakly convex WRS, bounded variance
DEPOSITUM [zhou2025decentralized] GT-M $\eta_{k}=\Theta(\frac{1}{\sqrt{K}}),k\in[K]$ Proximal, weakly convex WRS, bounded variance
DGF [swenson2022distributed] GD $\eta_{k}=\Theta(k^{-p}),p\in(0,1]$ Clarke regular, weakly coercive, decreasing consensus parameter WRS, bounded variance
ME-DOL [sahinoglu2024online] GD $\eta_{k}=\Theta(\frac{1}{\sqrt{K}}),k\in[K]$ Global Lipschitz continuous WRS, bounded variance, bounded second moment
@@
Our work (DSM) GD/GT /GD-M diminishing $\eta_{k}$ Path-differentiable, Coercive $R$
Table 1: A brief comparison of existing decentralized SGD-based methods for nonsmooth nonconvex optimization. Here, “GD”, “GT” and “GD-M”, ”GT-M” are abbreviations of “gradient descent”, “gradient tracking”, “gradient descent with momentum technique” and “gradient tracking with momentum technique”; “WRS” and “RR” are abbreviations of “with-replacement sampling” and “random reshuffling”; The term “conditionally controlled noise” means that the evaluation noise is bounded conditional on $\mathcal{F}_{k}=\sigma(\{{\bm{x}}_{i,l},l\leq k,i\in[d]\})$. The concept ”path-differentiable” is defined in Definition 2.6.
@@ 1.2 Existing works on stochastic subgradient-based methods for nonsmooth nonconvex optimization
In the training of nonsmooth neural networks, one major challenge comes from computing the (stochastic) subgradients of the loss function. In a wide range of deep learning packages, such as TensorFlow, PyTorch, and JAX, automatic differentiation (AD) algorithms have become the default choice for computing subdifferential. Utilizing the chain rule, AD algorithms recursively construct the so-called “subgradients” through the composition of Jacobians of each block of the neural network, yet they neglect the underlying nonsmoothness. As the chain rule fails for loss functions lacking Clarke regularity, the so-called “subgradient” may not necessarily be contained in its Clarke subdifferential (see examples in [bolte2020mathematical]). This limitation consequently renders the theoretical analyses of existing literature inapplicable to such functions.
@@
To characterize the behavior of AD algorithms, [bolte2021conservative] introduces the concept of path-differentiable functions, which constitute a broad function class, including semi-algebraic functions, semi-analytic functions, and functions whose graphs are definable in some o-minimal structures [davis2020stochastic]. More importantly, the set of path-differentiable functions is general enough to cover almost all loss functions in neural network tasks, encompassing most of the loss functions without Clarke regularity. The study [bolte2021conservative] further demonstrates that when applied to path-differentiable loss functions, the outputs of AD algorithms are contained within a conservative field, which is a generalization of the Clarke subdifferential. Conservative field admits chain rules for path-differentiable functions, thereby preserving several good properties. For a comprehensive discussion of this concept, readers are directed to Section 2.3.
@@
Based on the concepts of path-differentiable functions and conservative field, some recent works [bolte2021conservative, castera2021inertial, pauwels2021incremental, xiao2023adam, ruszczynski2020convergence, le2024nonsmooth, xiao2023convergence, bolte2022long] leverage the ordinary differential equation (ODE) approach [benaim2005stochastic, benaim2006dynamics, borkar2009stochastic, duchi2018stochastic] to study the behavior of stochastic subgradient-based methods in the non-distributed setting. Although these studies have achieved significant progress in establishing convergence properties for stochastic subgradient-based methods, the theoretical results of many approaches cannot avoid infinitely many spurious critical points due to the gap between the conservative field and the Clarke subdifferential [bianchi2022convergence, bolte2021nonsmooth]. Moreover, extending existing results to the multi-agent scenario is non-trivial, as the update schemes of all the agents are coupled through local communication under decentralized settings. Given the convergence properties that need improvement and limited existing results in the multi-agent scenario, developing a unified framework for establishing convergence properties for decentralized stochastic subgradient-based methods is of great importance and worth exploring.
@@ 1.3 A general framework for decentralized stochastic subgradient-based methods
${\bm{X}}_{k}$= $\left[{\bm{x}}_{1,k},{\bm{x}}_{2,k},\ldots,{\bm{x}}_{d,k}\right],$ (1.3)
$\Xi_{k+1}$= $\left[\xi_{1,k+1},\xi_{2,k+1},\ldots,\xi_{d,k+1}\right],$
${\bm{H}}_{k}$= $[\nabla f_{1}({\bm{x}}_{1,k}),\ldots,\nabla f_{d}({\bm{x}}_{d,k})].$
@@ 1.4 Contributions
We propose a unified framework (DSM) that encompasses a variety of decentralized stochastic subgradient-based methods. We verify that decentralized stochastic subgradient descent (DSGD), DSGD with gradient-tracking technique (DSGD-T) and DSGD with momentum (DSGD-M) all fit into the framework (DSM). In particular, a decentralized variant of SignSGD [bernstein2018signsgd], denoted as DSignSGD, is also enclosed by (DSM).
@@
Based on our theoretical results of the proposed framework (DSM), we prove that DSGD, DSGD-M, DSGD-T, and DSignSGD achieve global convergence in the random reshuffling case and high-probability global convergence in the with-replacement sampling case. We also show that with random initialization, DSGD, DSGD-M and DSignSGD can avoid the spurious critical points in the sense of conservative field in minimizing path-differentiable functions. Although these algorithms are widely used in the training of nonsmooth neural networks, existing theoretical results are limited to Clarke regular objective functions. Our theoretical findings first fill the gap in solving Problem (DOP) with objectives lacking Clarke regularity.
@@ 2.1 Notations
We denote $\langle\cdot,\cdot\rangle$ as the standard inner product, while $\|\cdot\|$ as the $\ell_{2}$-norm of a vector or spectral norm of a matrix. We refer to $\|\cdot\|_{1}$ as $\ell_{1}$-norm of a vector and $\|\cdot\|_{\mathrm{F}}$ as Frobenius norm of a matrix. $\mathbb{B}({\bm{x}},\delta):=\left\{\tilde{{\bm{x}}}\in\mathbb{R}^{n}:\|\tilde{{\bm{x}}}-{\bm{x}}\|^{2}\leq\delta^{2}\right\}$ refers to the ball centered at $\bm{x}$} with radius $\delta$. For a given set $\mathcal{Y},\operatorname{dist}({\bm{x}},\mathcal{Y})$ denotes the distance between $\bm{x}$} and a set $\mathcal{Y}$, i.e. $\operatorname{dist}({\bm{x}},\mathcal{Y}):=\arg\min_{y\in\mathcal{Y}}\|{\bm{x}}-{\bm{y}}\|,$ $\mathrm{conv}\,\mathcal{Y}$ denotes the convex hull of $\mathcal{Y}$, and $\mathcal{Y}^{d}$ denotes $d$-fold Cartesian product. The notation $\otimes$ stands for the Kronecker product. $\Delta_{m}:=\{(\lambda_{0},\ldots,\lambda_{m}):\lambda_{i}\geq 0,\sum_{i=0}^{m}\lambda_{i}=1\}$ stands for the simplex of dimension $m$.
@@
Moreover, we denote $\mathcal{N}_{i}$ as the set of neighbor agents of agent $i$ together with $\{i\}$, $\mathbb{R}_{+}$ as the set of all nonnegative real numbers. For any $N>0,[N]:=\{1,\ldots,N\}$. And let ${\bm{1}}_{d}\in\mathbb{R}^{d}$ represent a vector of all $1$’s, ${\bm{e}}_{i}\in\mathbb{R}^{d}$ represent $0,\ldots,1,\ldots,0]^{\top}$, where $1$ is the $i$-th component. For two integers $i$ and $j$, $i\land j$ represents $\min\{i,j\}$.